OpenStudy (anonymous):

8/9a^2 divide 4a^2-4a-24/a^2-6a+9 Perform the indicated operation.

OpenStudy (anonymous):

@coral

OpenStudy (anonymous):

i ned help please @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

\[\LARGE{\frac{8}{9a^2} \div \frac{4a^2-4a-24}{a^2-6a+9}}\]\[\LARGE{\frac{8}{9a^2} \div \frac{4a^2-12a+8a-24}{a^2-3a-3a+9}}\]\[\LARGE{\frac{8}{9a^2} \div \frac{ 4a(a-3)+8(a-3)}{a(a-3)-3(a-3)}}\]\[\LARGE{\frac{8}{9a^2} \div \frac{(4a+8) \cancel {(a-3)}}{(a-3) \cancel {(a-3)}}}\]\[\LARGE{\frac{8}{9a^2} \times \frac{(a-3)}{(4a+8)}}\]